### v1

@@ 2.2 Medium-range weather forecasting
Overall, XiChen achieves a skillful weather forecasting lead time of over 8.25 days (with ACC of Z500 $>0.6$), thereby achieving performance comparable to GFS (8.25 days) while substantially outperforming both Aardvark (less than 8 days) [22] and GraphDOP (5 days) [23]. This achievement is particularly noteworthy considering that XiChen was trained using only 12 years of ERA5 data at a 1.40625 resolution. Moreover, XiChen assimilated a limited subset of observations used in operational NWP systems–including GDAS prepbufr, AMSU-A, and MHS observations–significantly fewer than those used by GFS, Aardvark [22], and GraphDOP [23]. XiChen’s RMSE surpasses that of IFS HRES at larger lead times primarily because its forecasts tend to favor smoothing, as evidenced by the energy spectra presented in Figure S8. This tendency is a common characteristic observed in contemporary AI-based weather forecasting models [8, 9, 10]. When initialized using ERA5 data, XiChen demonstrates a skillful forecast lead time of more than 10 days. Further details can be found in the Supplementary Text and Figure S5.
2.3 Contribution of assimilating satellites
As illustrated in Figure S1, XiChen employs a cascade DA framework, which provides the flexibility to select specific observations for assimilation to evaluate the impact of individual satellite observations on the performance of the analysis field. Notably, since the GDAS prepbufr data includes observations for all variables, our approach begins by assimilating GDAS prepbufr observations to effectively estimate the entire atmospheric field. Subsequently, satellite observations are assimilated to further refine the corresponding key variables.
[…]
Figure 3(a) illustrates the impact of separately assimilating AMSU-A or MHS, as well as simultaneously assimilating these two satellite observations, on the analysis quality of upper-air variables, based on the assimilation of GDAS prepbufr data. The first column in each scorecard represents the RMSEs of the analysis field when only GDAS prepbufr data is assimilated. The second column shows the results of additional assimilation of AMSU-A, while the third column reflects the impact of assimilating MHS. The fourth column represents the outcome of assimilating AMSU-A followed by MHS, and the fifth column shows the results of assimilating MHS followed by AMSU-A.
The assimilation of AMSU-A channels 5 to 10, which are primarily designed to observe atmospheric temperature profiles [39], directly improved the upper-air temperature fields. Furthermore, the assimilation of AMSU-A data effectively corrected the upper-air geopotential fields and positively influenced the wind fields. This outcome is likely due to changes in the temperature implicitly adjusting the pressure gradient force, which subsequently modifies the geopotential and wind fields. These findings suggest that improvements in the temperature propagate to the geopotential through their dynamic relationship.
Similarly, the assimilation of channels 3 to 5 of the MHS, intended primarily for observing atmospheric humidity profiles [39], directly enhances the upper-air specific humidity. However, this process adversely affects the upper-air temperature and geopotential fields. This is likely because MHS brightness temperature observations are influenced by factors such as clouds, precipitation, and surface radiation, which are not currently accounted for by XiChen. Furthermore, the extensive adjustments in specific humidity introduced during the DA process may create transient imbalances that disrupt the dynamical state of the atmosphere. Such imbalances often manifest as gravity waves or other types of dynamical feedback, causing the analysis states of the geopotential and wind fields to diverge from equilibrium and thereby amplifying errors. In the future, the incorporation of physical equilibrium constraints into XiChen may be considered as a potential solution to this problem.
Nevertheless, the simultaneous assimilation of AMSU-A and MHS effectively coordinates upper-air atmospheric fields, reducing analysis errors caused by the sparse upper-air observations in GDAS prepbufr. Additionally, the joint assimilation of AMSU-A and MHS mitigates the adverse effects of MHS-only assimilation on the atmospheric temperature and geopotential fields. Importantly, changes in the assimilation order of AMSU-A and MHS have minimal impact on the quality of the analysis fields, highlighting the robustness and scalability of XiChen.
Figure 3(b) presents the medium-range weather forecast scorecard comparing the additional assimilation of two satellite observations, AMSU-A and MHS, to the baseline that uses only GDAS prepbufr observations. The scorecard displays the RMSE values for the corresponding variables at specific pressure levels. In terms of RMSE, the additional assimilation of AMSU-A and MHS demonstrates improved performance across nearly all lead times for upper-air variables, particularly for geopotential, temperature, and specific humidity. These results align with findings from the DA cycle experiment, suggesting that incorporating these satellite observations effectively enhances the analysis of corresponding variables and positively impacts subsequent forecasts. At the pressure level of 50 hPa, forecasts incorporating both AMSU-A and MHS data show significant improvement. This enhancement is largely due to the sparsity of GDAS prepbufr conventional observations in the upper atmosphere. Satellite observations effectively address this data gap, improving medium-range weather forecasts in upper atmospheric levels. For results on the effect of additional assimilation only of AMSU-A or MHS on the weather forecasts after assimilating GDAS prepbufr, see Figure S6 and Figure S7.
@@ 2.4 Contribution of the 4DVar knowledge
We performed a one-year DA cycle experiment, with results presented in Figure 4. The findings indicate that as the DAW length increases, the analysis error decreases, highlighting the positive impact of including more temporal observations on XiChen’s performance. Notably, even without the 4DVar knowledge constraint (DAW [0h]), XiChen maintained a stable DA cycle over one year, despite producing larger analysis errors. This result further demonstrates the robustness of the XiChen model. Additionally, we conducted a 10-day medium-range weather forecast experiment, with results provided in Figure S9. Forecasts generated from analysis fields with longer DAWs exhibited lower RMSE and higher ACC, further confirming that incorporating longer temporal information through 4DVar knowledge enhances forecasting accuracy. These results collectively demonstrate the effective improvement in analysis and forecasting performance achieved by introducing 4DVar knowledge, thereby pushing the fully AI-driven global weather forecasting system towards operationalization.
@@ 2.5 Assimilation and forecasting of tropical cyclones
To evaluate the track performance of TCs using XiChen, the IBTrACS is employed as the reference dataset. IBTrACS, accessible at here, consolidates recent and historical TC data from multiple agencies, providing the most comprehensive global collection of TC information.
To assess XiChen’s ability to assimilate and forecast TC trajectories, we employ the TempestExtremes tracker [37] to identify TC tracks. For comparison with the IBTrACS dataset and ERA5 reanalysis, we adopt the configuration used in NeuralGCM [38], with parameter details listed in Table S7. To maintain consistency with IBTrACS, we exclude any TCs identified by this configuration that are not present in the IBTrACS dataset, ensuring that only IBTrACS TCs are included in our results. For details, please see the Supplementary Text.
@@
Figure 5: Assimilation and Forecasting of Tropical Cyclones. (a) Tropical cyclone (TC) trajectory tracks from IBTrACS ground truth data for the entire year of 2023. (b) TC trajectory tracks (blue lines) from the ERA5 reanalysis for 2023. (c) TC trajectory tracks (blue lines) from the XiChen analysis fields obtained by assimilating GDAS prepbufr, AMSU-A, and MHS during a one-year data assimilation (DA) cycle. (d) TC trajectory tracks (blue lines) from the XiChen analysis fields obtained by assimilating GDAS prepbufr and MHS during a one-year DA cycle. (e) TC trajectory tracks (blue lines) from the XiChen analysis fields obtained by assimilating GDAS prepbufr and AMSU-A during a one-year DA cycle. (f) TC trajectory tracks (blue lines) from the XiChen analysis fields obtained by assimilating only GDAS prepbufr during a one-year DA cycle. In panels (b) to (f), the IBTrACS ground truth tracks are shown in gray. (g) Comparison of XiChen performance in terms of mean direct position error over 5-day TC forecasts in 2023. (h) Comparison of XiChen performance in terms of average maximum wind speed error over 5-day TC forecasts in 2023. (i) Comparison of XiChen performance in terms of average minimum MSLP error over 5-day TC forecasts in 2023. In panels (g)-(i), blue lines represent results initialized by assimilating GDAS prepbufr only, while yellow, green, and red lines represent results from additionally assimilating AMSU-A, MHS, and both AMSU-A and MHS, respectively.
@@
This study evaluates XiChen’s performance in assimilating and forecasting the trajectories of TCs. Figure 5(a)-(f) provides a comparative analysis of TC tracks from different datasets and XiChen’s analysis fields. Figure 5(a) depicts the TC trajectories from the IBTrACS ground truth dataset for 2023, including 81 TCs. The blue lines in Figure 5(b) depict the TC trajectories identified in the ERA5 reanalysis dataset for 2023 that correspond with TC trajectories in IBTrack, comprising a total of 52 tropical cyclones. The ERA5 dataset successfully identified 64.2% of the TCs documented in IBTrack, with an average trajectory deviation of 0.66 degrees. The blue lines in Figures 5(c)-(f) represent TC trajectories generated by XiChen during the one-year DA cycle. These trajectories were obtained through various assimilation configurations: GDAS prepbufr with AMSU-A and MHS; GDAS prepbufr with MHS; GDAS prepbufr with AMSU-A; or GDAS prepbufr alone. XiChen identified one more TC following the assimilation of AMSU-A data compared to the assimilation of only GDAS prepbufr data. This improvement is attributed to the enhanced temperature field resulting from AMSU-A assimilation, which subsequently enhances the geopotential field. In turn, this indirectly improves the wind field through a dynamic equilibrium relationship. A more accurate wind field analysis facilitates the identification of closed circulations associated with low-pressure centers, thereby allowing for more precise identification of TCs. Conversely, the number of TCs identified through the additional assimilation of MHS decreased due to its negative impact on the temperature and wind fields. Although the number of identified TCs varies slightly across different observation assimilation scenarios, the overall TC distribution patterns remain consistent, with comparable average trajectory errors. This consistency demonstrates XiChen’s robust performance. Notably, the TC trajectories generated by XiChen include some TCs not identified in ERA5, suggesting that XiChen not only reproduces ERA5 data but also exhibits a degree of generalization capability. Figures S10 through S13 display the results of tropical cyclone trajectory identification from forecasting fields across different forecast lead times.
Furthermore, to assess trajectory and intensity errors in XiChen’s TC forecasts, we initiated 5-day forecasts from the start time of all TCs documented in the IBTrack dataset. The tracking and error calculations were performed at 6-hour intervals. The methodology for tracking TCs in the forecasting fields followed the approach established in the Pangu-Weather [5]. Figures 5(g)–(i) compare four assimilation configurations: (1) assimilating GDAS prepbufr only (blue), (2) assimilating GDAS prepbufr and AMSU‑A (yellow), (3) assimilating GDAS prepbufr and MHS (green), and (4) assimilating GDAS prepbufr, AMSU‑A, and MHS (red). Figure 5(g) shows that assimilating MHS alone increases track errors beyond 48 h, while additional assimilating AMSU‑A mitigates this deterioration. The assimilation of GDAS prepbufr and AMSU‑A does not yield a track improvement relative to GDAS prepbufr only. In Figures 5(h) and 5(i), both maximum wind speed and minimum MSLP errors decrease substantially after assimilating MHS, likely due to improved water‑vapor correction in the atmospheric state. Assimilating AMSU‑A similarly enhances intensity forecasts, and the combined assimilation of AMSU‑A and MHS achieves the greatest reduction in intensity errors. The intensity forecasts after 72 hours exhibit improved accuracy, likely because it is produced by XiChen‑Medium. XiChen‑Medium is fine-tuned using XiChen‑Short’s 5‑day forecast fields as initial inputs, thereby reducing long‑term forecast errors. For details on XiChen‑Short and XiChen‑Medium, see Section 4.4. Notably, our MSLP errors across all lead times closely match those of FuXi-Extreme [40], a medium-range forecasting and post-processing model trained at 0.25 resolution, demonstrating the strong potential of our model for operational use.
3 Disscusion

### v2

@@ 2.2 Medium-range weather forecasting
When using XiChen’s self-generated initial conditions, XiChen achieved a skillful weather forecasting lead time exceeding 8.75 days, surpassing the GFS system (8.25 days). Notably, for key variables such as T2M, T850, MSLP, U10, and V10, the fully AI-driven global weather forecasts of XiChen deliver lower RMSEs compared to the GFS for nearly the entire 10-day lead time. The RMSEs for all variables approach or even surpass the performance of the IFS HRES as the lead time increases. For Q500, XiChen exceeds the GFS after 2 days and the IFS HRES after 4 days. Although the forecast error for Z500 remains higher than that of both the GFS and the IFS HRES during the first 6 days, XiChen surpasses the GFS after 7 days and outperforms IFS HRES after 8 days. This achievement is particularly noteworthy considering that XiChen was trained using only 12 years of ERA5 data at a 1.40625 resolution. Moreover, XiChen assimilated a limited subset of observations used in operational NWP systems–including GDAS prepbufr, MHS, ASCAT, AMSU-A, and SATWND observations–significantly fewer than those used by GFS and IFS HRES. XiChen’s RMSE surpasses that of IFS HRES at larger lead times, which may be because its forecasts tend to favor smoothing, as evidenced by the energy spectra presented in Figure S5. This tendency is a common characteristic observed in contemporary AI-based weather forecasting models [lam2023learning, ailabchen2025fengwu, fudanchen2023fuxi]. For comprehensive insights into the 10-day medium-range forecast RMSE scorecards across all variables for Xichen versus GFS, and XiChen versus IFS HRES, kindly consult the Supplementary Text, as well as Figures S6 and S7.
2.3 Contribution of assimilating different observations
As illustrated in Figure S1, a pivotal innovation of XiChen lies in its scalable capability to integrate multi-modal observational data within a cascade DA framework for the precise estimation of initial conditions. With the continuous expansion in both the number and diversity of observational modalities, a pressing question emerges: which specific observations exert the most significant influence on medium-range weather forecasting accuracy [laloyaux2025using]? To address this inquiry, we designed an ablation experiment aimed at quantifying the marginal impact of omitting individual observation sources on 10-day medium-range forecast errors.
In this experiment, various observation sources were systematically excluded from the assimilated dataset, after which DA was conducted using XiChen. The resulting 10-day forecast mean errors were subsequently compared against those derived from assimilating the complete set of available observations (designated as the “All” configuration). For instance, configurations such as “No ASCAT” and “No Satellites” denote the exclusion of ASCAT data and all satellite-derived observations, respectively, from the “All” configuration. Figure 3 provides a detailed depiction of the percentage increase in the 10-day mean latitude-weighted RMSE across all test cases, benchmarked against the “All” configuration.
The results reveal that removing any observation source increases errors in medium-range global weather forecasts. Among these, GDAS prepbufr has the most significant impact, likely due to its direct measurements of fundamental atmospheric variables. Furthermore, removing ASCAT or SATWND increases errors across all variables. Since ASCAT and SATWND observe surface winds and upper-air winds, respectively, this highlights XiChen’s effectiveness in learning correlations among atmospheric variables. This capability enables the system to use observed variables to correct unobserved ones.
Figure 3: Headline scorecards of the average normalized RMSEs for 10-day medium-range weather forecasting, based on the assimilation of different observations. The colors denote the percentage difference relative to the baseline, which assimilates all observations, including GDAS prepbufr, AMSU-A, MHS, ASCAT, and SATWND. Blue shading indicates a reduction in normalized RMSE, while red shading indicates an increase.
Removing data from channels 5 to 10 of the AMSU-A satellite, which are primarily used to observe atmospheric temperature profiles, resulted in increased errors in upper-air temperatures [wang2014amsu]. This suggests that the observation operator in XiChen effectively captures AMSU-A’s sensitivity to atmospheric temperature profiles. Moreover, excluding AMSU-A observations increased errors in the upper-air geopotential fields and introduced errors in the wind fields. These effects likely arise from XiChen’s ability to implicitly adjust the pressure gradient force in response to temperature variations, thereby influencing the potential and wind fields. Additionally, the exclusion of AMSU-A data led to increased humidity errors, highlighting XiChen’s ability to learn the temperature-humidity balance relationship.
Similarly, the exclusion of data from channels 3 to 5 of the MHS, which are predominantly utilized for observing atmospheric humidity profiles [zou2017impacts], led to increased errors in upper-air specific humidity. This finding underscores the effectiveness of the observation operator within the XiChen framework in accurately encapsulating MHS’s sensitivity to atmospheric humidity profiles. Moreover, the omission of MHS observations was observed to introduce elevated errors across other variables to a certain degree. This observation further emphasizes XiChen’s capability to leverage its learned equilibrium constraints, effectively redistributing the specific humidity adjustments to other related variables within the system.
Figure 4: Headline scorecards of the global-averaged latitude-weighted RMSEs for 10-day medium-range weather forecasting, based on the assimilation of different observations. The 10-day medium-range forecasts based on analysis fields that assimilate all observational data—including GDAS prepbufr, AMSU-A, MHS, ASCAT, and SATWND—are compared with the forecasts based on analysis fields that assimilate only the GDAS prepbufr. The colors indicate the percentage difference relative to the baseline, which assimilates only the GDAS prepbufr. Blue shading represents a reduction in RMSE, while red shading represents an increase.
Notably, the exclusion of all satellite observations results in a substantially greater increase in forecast errors compared to the removal of a single satellite observation type. This highlights the robust capability of XiChen to effectively integrate a diverse array of satellite observations, thereby enhancing forecast accuracy. Additionally, Figure 4 illustrates the influence of additional assimilating data from MHS, ASCAT, AMSU-A, and SATWND, alongside the GDAS prepbufr, on the RMSE of 10-day medium-range weather forecasts. Across nearly the entire forecast horizon, the incorporation of satellite observations markedly reduces forecast errors for all variables. This reduction is particularly pronounced for upper-atmospheric variables, such as geopotential height, temperature, and specific humidity, underscoring the critical role of satellite data in improving the accuracy of medium-range weather predictions.
2.4 Physical Consistency of the Assimilation
This section examines how XiChen account for the impacts of assimilating a single perturbation observation on the background field and evaluates whether this process adheres to fundamental physical principles.
In this experiment, two separate DA runs were conducted. The first run used ERA5 as the background field to assimilate raw satellite observations. In the second run, a perturbation is introduced to the raw satellite inputs across all observation channels individually. The differences between the two DA results illustrate the impact of the perturbation on the analysis fields. The DA time was set to 00 UTC on July 24, 2023, and perturbation experiments were conducted separately for AMSU-A and MHS. The perturbation location, selected at 20 N, 120 E, was near the path of Typhoon Doksuri during this period. The perturbation magnitude was set to +5 K, and independent experiments were performed for each satellite channel.
Figures 5 and 6 present the horizontal and vertical distributions of analysis field variations resulting from individually perturbing three AMSU-A channels (5, 6, and 7) and three MHS channels (3, 4, and 5), respectively. For AMSU-A, increases in brightness temperature correspond to rising atmospheric temperatures, consistent with radiative transfer theory, demonstrating that XiChen has effectively learned AMSU-A’s temperature sensitivity. For MHS, increases in brightness temperature correspond to decreases in specific humidity, indicating that XiChen has similarly captured MHS’s sensitivity to atmospheric moisture. In both figures, the analysis increments exhibit propagating variations horizontally and vertically and display localized, anisotropic structures around the observations that align with the background geopotential field, reflecting the flow-dependent nature of the XiChen system. Additionally, for AMSU-A, anticyclonic wind responses surrounding temperature increments imply that XiChen adheres to certain equilibrium constraints. Overall, these results are consistent with findings from 4DEnVar [zhu2022four].
Figure 5: Changes in the analysis fields resulting from a 5 K perturbation to the AMSU-A observation at the selected location, where the background field is the ERA5 data at 00 UTC on July 24, 2023. The perturbation was introduced at 03 UTC, three hours after the analysis time, near Typhoon Doksuri at 20 N, and 120 E (indicated by the yellow dot). The first row displays the horizontal spatial distribution of analysis changes for channels 5 to 7 at 700, 200, and 200 hPa. Shading represents temperature differences (units: $K$), while vectors indicate wind differences (units: $ms^{-1}$) resulting from the assimilation of single-point perturbation observations. The solid contour illustrates the background geopotential field (units: $m^{2}s^{-2}$). The second row displays the vertical distribution along the same west-east cross-section.
Figure 6: Changes in the analysis fields resulting from a 5 K perturbation to the MHS observation at the selected location, where the background field is the ERA5 data at 00 UTC on July 24, 2023. The perturbation was introduced at 03 UTC, three hours after the analysis time, near Typhoon Doksuri at 20 N, and 120 E (indicated by the yellow dot). The first row shows the horizontal spatial distribution of analysis changes for channels 3 to 5 at 300, 500, and 500 hPa. Shading represents differences in specific humidity (units: $gkg^{-1}$), while vectors indicate wind differences (units: $ms^{-1}$) resulting from the assimilation of single-point perturbation observations. The solid contour illustrates the background geopotential field (units: $m^{2}s^{-2}$). The second row displays the vertical distribution along the same west-east cross-section.
@@ 2.5 Contribution of the 4DVar knowledge
We performed a one-year DA cycle experiment, with results presented in Figure 7. The findings indicate that as the DAW length increases, the analysis error decreases, highlighting the positive impact of including more temporal observations on XiChen’s performance. Notably, even without the 4DVar knowledge constraint (DAW [0h]), XiChen maintained a stable DA cycle over one year, despite producing larger analysis errors. This result further demonstrates the robustness of the XiChen model. Additionally, we conducted a 10-day medium-range weather forecast experiment, with results provided in Figure S8. Forecasts generated from analysis fields with longer DAWs exhibited lower RMSE and higher ACC, further confirming that incorporating longer temporal information through 4DVar knowledge enhances forecasting accuracy. These results collectively demonstrate the effective improvement in analysis and forecasting performance achieved by introducing 4DVar knowledge.
@@ 2.6 Assimilation and forecasting of tropical cyclones
Figure 8: Assimilation and Forecasting of Tropical Cyclones. (a) Tropical cyclone (TC) trajectory tracks from IBTrACS ground truth data for the entire year of 2023. (b) TC trajectory tracks (blue lines) from the ERA5 reanalysis for 2023. (c) TC trajectory tracks (blue lines) from the XiChen analysis fields obtained by assimilating GDAS prepbufr, AMSU-A, and MHS during a one-year data assimilation (DA) cycle. (d) Comparison of XiChen performance in terms of mean direct position error over 5-day TC forecasts in 2023. (e) Comparison of XiChen performance in terms of average maximum wind speed error over 5-day TC forecasts in 2023. (f) Comparison of XiChen performance in terms of average minimum MSLP error over 5-day TC forecasts in 2023. In panels (d) through (f), the blue lines denote the outcomes initialized with ERA5 data, whereas the yellow and red lines correspond to results derived from analysis fields obtained by assimilating GDAS prepbufr exclusively and by further incorporating all satellite observations, respectively.
To assess XiChen’s ability to assimilate and forecast TC trajectories, we employ the TempestExtremes tracker [ullrich2021tempestextremes] to identify TC tracks. For comparison with the IBTrACS dataset and ERA5 reanalysis, we adopt the configuration used in NeuralGCM [kochkov2024neural], with parameter details listed in Table S9. To maintain consistency with IBTrACS, we exclude any TCs identified by this configuration that are not present in the IBTrACS dataset, ensuring that only IBTrACS TCs are included in our results. For details, please see the Supplementary Text.
@@
This study evaluates XiChen’s performance in assimilating and forecasting TC trajectories. Figures 8(a)-(c) provide a comparative analysis of TC tracks from different datasets and XiChen’s analysis fields. Figure 8(a) depicts the TC trajectories from the IBTrACS ground truth dataset for 2023, including 81 TCs. The blue lines in Figure 8(b) depict the TC trajectories identified in the ERA5 reanalysis dataset for 2023 that correspond with TC trajectories in IBTrACS, comprising a total of 52 tropical cyclones. The ERA5 dataset successfully identified 64.2% of the TCs documented in IBTrACS, with an average trajectory deviation of 0.66 degrees. The blue lines in Figure 8(c) represent TC trajectories generated by XiChen during the one-year DA cycle. The XiChen analysis fields successfully identified 61.73% of the TCs documented in IBTrACS, with an average trajectory deviation of 0.79 degrees. These results demonstrate that XiChen effectively leverages multi-source observational data to retain TC information within the analysis field. Notably, the TC trajectories generated by XiChen include some TCs not identified in ERA5, suggesting that XiChen not only reproduces ERA5 data but also exhibits a degree of generalization capability. Figure S9 displays the results of TC trajectory identification from forecasting fields across different forecast lead times.
To quantitatively evaluate the impact of satellite observations on track and intensity errors in XiChen‘s TC forecasts, we conducted 5-day predictions starting from the onset times of all typhoons in the IBTrACS dataset. Tracking and error calculations were performed at 6-hour intervals. The typhoon tracking methodology follows the scheme established by the Pangu-Weather system [bi2023accurate]. Figures 8(d)–(f) compare forecasts initiated with three different initial fields: (1) ERA5 data (blue), (2) assimilation of only GDAS prepbufr data (yellow), and (3) simultaneous assimilation of GDAS prepbufr, MHS, ASCAT, AMSU-A, and SATWND data (red). Figure 8(d) shows that incorporating satellite observations into GDAS prepbufr assimilation significantly improves the accuracy of TC track forecasts. However, since the forecast model was trained using ERA5 data, its performance when using ERA5 as the initial field remains superior. In Figures 8(e) and 8(f), XiChen demonstrates similar maximum wind speed errors and minimum mean sea level pressure errors across all three initial fields. This is primarily due to biases in the ERA5 dataset regarding TC intensity, which were learned by XiChen during training.
3 Discussion